OpenStudy (anonymous):

write a quotient of two cube roots so that the answer, when simplified is 3a^2 over 4b63

OpenStudy (anonymous):

Do this: \[ \frac{3a^2}{4b63}=\frac{\sqrt[3]{(3a^2)^3}}{\sqrt[3]{(4b63)^3}} \]

OpenStudy (anonymous):

the 6 isnt suppose to be there ha